OpenStudy (anonymous):

Choose the correct simplification of the expression (3xy4)2(y2)3. 6x2y14 9x2y14 9x3y11 6x3y11

OpenStudy (anonymous):

Please use the Equation editor, the problem is not clear. Does it look like this? \[(3x * y^4) 2 (y^2)^3\]

Nnesha (nnesha):

it's \[\huge\rm (3xy^4)^2(y^2)^3\]

OpenStudy (anonymous):

Yes, like @Nnesha did it :]

Nnesha (nnesha):

apply the exponent rules \[\huge\rm (x^m)^n=x^{m \times n}\]

Nnesha (nnesha):

and \[\huge\rm (ab)^x =a^x b^x\]

OpenStudy (anonymous):

Oh okay! So \[y^8? correct?\]

Nnesha (nnesha):

hmm okay so you are saying (3xy^4)^2 = y^8 ?

OpenStudy (anonymous):

Oh no. I am confused (´・_・`)

Nnesha (nnesha):

\[\huge\rm (ab c^m)^x =a^x b^x c^{m \times x}\] (3xy^4)^2 should be like this

Nnesha (nnesha):

2 is an exponent of all numbers and variables in the parentheses (3xy^4)^2 = ?

OpenStudy (anonymous):